• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name slifer

    Log entry time 08:07:36 on October 13,1998

    Entry number 2783

    keyword=shift summary

    Finished up Nitrogen Runs
    and Empty cell runs (Coll in /Coll out).
    Called for access to reset Q3 which had quenched
    He-3 pressure curve runs delayed as we had concern
    about shower/preshower phototube readouts.

    Last NMR measurement of "Be Happy" gives ~14% at 07:10